OpenStudy (anonymous):

Question of utmost urgency, 4 hours for my exam!!!!! Finding all the points on the curve at which the tangent passes through the origin y = 4x^3 - 2x^5 How do I approach and solve this?

OpenStudy (anonymous):

@phi @zepdrix

OpenStudy (anonymous):

@WordGEEK

OpenStudy (anonymous):

Sorry I do not know this :(

OpenStudy (anonymous):

Do you know anyone who does? I really need to learn this :(

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

@ganeshie8

OpenStudy (kainui):

Not gonna lie, this is a tough one. It seems like graphing this would be one of the easiest ways to see, then take a line through the origin and just rotate it. When it first hits the graph, that's exactly where the tangent line would fall. Interesting question, I'm sure there's a better way to do this, let me think about it a little.

OpenStudy (kainui):

So the general lines we're looking to compare this to are of the form: y=mx since the y-intercept is 0. We also know that because m=rise/run and that this line goes through the origin: \[\Large m= \frac{y_2-y_1}{x_2-x_1}=\frac{y_2}{x_2}\]because (x1,y1)=0. That's actually not so bad now that we notice this. Notice all I did here wasn't any kind of magic genius, all I did was begin to start looking at the definitions of things and plugging in numbers and it's getting us closer to the answer.

OpenStudy (kainui):

(x1, y1)=(0,0) is what i meant to write haha. Ok, let's take this a little further. The slope of the line we're making is just the y value divided by the x value at that point. This means that for a line to have the same slope as our line, we just compare their slopes! \[\Large \frac{y}{x}=\frac{dy}{dx}\] Now we just solve this formula by plugging in. Does that make sense? y/x is just the slope of our linear line through the origin and dy/dx is the slope of the curve we were given. When these are equal, we will have a tangent line at the same place on our curve because we set them equal to each other.

OpenStudy (anonymous):

so we're actually using y - y_1 = m(x - x_1) y - 0 = m(x - 0) so its y=mx What we do next is find f'(x)? f'(x) = 12x^2 - 10x^4?

OpenStudy (kainui):

Yes exactly, now we're just setting the derivative equal to y/x since that's the slope of our line.

OpenStudy (anonymous):

Okay our next step is to put our f'(x) into y = mx?

OpenStudy (kainui):

It actually solves very nicely! To phrase it in another notation, but identical to what I have above:\[\Large f'(x)=\frac{f(x)}{x}\] I have to go eat dinner right now I'm afraid. I'll be back very soon to help you out though, I want to help make sure you understand so you can do well! =)

OpenStudy (anonymous):

I'll have to go by then.. I'm leaving right now as my exam center is over 10 km away :( I'm nervous as hell..

OpenStudy (kainui):

Good luck! Try not to freak out too much and breathe.

OpenStudy (anonymous):

Thanks bud :) I'll try my best! bye! :) I'll be back this afternoon